OpenStudy (anonymous):

PLZ HELP-WILL FAN Neil has been running a tutoring business since 2005. He charges a monthly fee for weekly tutoring sessions and a phone help line. Each year, he has increased his fee by the same amount. The table shows what Neil charged each customer for two given years of his business: Year Annual Tutoring Fee 2005 $1200 2008 $1350 A. What is the rate of change and initial value for Neil’s business? How do you know? B. Write an equation in slope-intercept form to represent the fees that Neil charges each year.

OpenStudy (anonymous):

@Nnesha @kevin_bridges @CandiceKing @saseal

OpenStudy (anonymous):

A: is a change of 150 every 3 years and I don't know B

OpenStudy (anonymous):

@Ashton_17 @ashleyz

OpenStudy (anonymous):

@Xep

OpenStudy (anonymous):

\[1350-1200=150\]\[2008-2005=3\]\[\frac{ 150 }{ 3 } =50\]

OpenStudy (anonymous):

@tiffanymak1996

OpenStudy (anonymous):

@saseal it came up like this\[1350-1200=150\]\[2008-2005=3\]\[\frac{ 150 }{ 3 } =50\]

OpenStudy (anonymous):

@Anthony2992992

OpenStudy (anonymous):

\[y=50x+1200\]assuming 2005 is year 0

OpenStudy (anonymous):

but what do i write for part b

OpenStudy (anonymous):

that one up there

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

B. Write an equation in slope-intercept form to represent the fees that Neil charges each year.

OpenStudy (anonymous):

I'm really confused what to write @saseal

OpenStudy (anonymous):

since he started business in 2005 so year 0 is x=0

OpenStudy (anonymous):

thats the equation is slope intercept form?

OpenStudy (anonymous):

when year 0 his price is 1200 that is the y-intercept

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

x=0?

OpenStudy (anonymous):

x=0 then y=1200

OpenStudy (anonymous):

try x=3 see if you get 1350

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

with?

OpenStudy (anonymous):

how to get the equation

OpenStudy (anonymous):

you learned y=mx+c ?

OpenStudy (anonymous):

i did y=mx+b

OpenStudy (anonymous):

thats the equation for slope right?

OpenStudy (anonymous):

its the same thing

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thats what they want you to do

OpenStudy (anonymous):

so whats the final answer?

OpenStudy (anonymous):

@saseal ^

OpenStudy (anonymous):

is it y=50x+3? @saseal

OpenStudy (anonymous):

y=50x+1200

OpenStudy (anonymous):

can u help me with another? @saseal